What is the y-intercept for the graph of 5x + 2y = –8?
Answer by Cromlix(4381)

Rearrange 5x + 2y = -8
into y = mx + c form.
5x + 2y = -8
2y = -5x - 8
y = -5/2x - 4
y-intercept is (0, -4)
Hope this helps.
:-)